# A Survey on Diffusion Language Models

### v1

@@ 2.4 Hybrid AR-Diffusion Language Models
Hybrid AR-Diffusion models aim to strike a balance between the full parallelism of non-autoregressive models and the strong causal dependency modeling of autoregressive models. A prominent strategy for hybrid AR-diffusion modeling adopts a block-wise semi-autoregressive generation process. In this setting, the model generates blocks of tokens autoregressively, while the tokens within each block are generated in parallel using a diffusion-like iterative process. Early efforts such as SSD-LM [74] pioneered hybrid approaches by a block-wise continuous diffusion process on simplex representations, AR-DIFFUSION [75] illustrates a multi-level diffusion process and achieves semi-autoregressive by adjusting timestep according token position. Recent Representative model BD3-LM [76] further advances this direction on discrete models, demonstrating strong performance compared to pure AR and diffusion models. CtrlDiff [77] improves this paradigm by introducing dynamic block prediction techniques to enhance block-level efficiency and control.
@@
Beyond block-based approaches that combine AR and diffusion at sequence level, hybridization can also occur at the architectural level, where some part of the neural network, typically the encoder, diffuses the entire sequence altogether to an intermediate representation, then an autoregressive decoder generates the final sequence [143]. LADIDA [144] is a slightly different approach that diffuses at document level but decodes sentences by an AR decoder. SpecDiff [78] proposed a collaborative speculative decoding framework, where a lightweight diffusion model drafts candidate outputs, which are then validated and finalized by a large AR model.
@@ 3.1 Pre-training and Supervised Fine-tuning
To accelerate training, particularly for large-scale models, it is common practice to initialize DLMs from pretrained AR language models or image diffusion models. DiffuGPT and DiffuLLaMA [27] try to initialize masked DLMs with open-sourced LLMs from 127M to 7B parameters, found that DLMs can be efficiently adapted from AR models, significantly reducing training time and cost while achieving comparable or even superior performance to their AR counterparts. Building on this insight, Dream-7B is initialized from Qwen 2.5 7B [142], and is reported to outperform both LLaDA-8B and LLaMA3-8B on various benchmarks. Some multimodal DLMs, on the other hand, are initialized from pretrained image diffusion models. D-DiT [79] and Muddit [80] are initialized from pretrained MM-DiT backbones from SD3 [11] and Meissonic [145] respectively. Although these models are not originally designed for text generation, their latent representations contain intrinsic language-aligned knowledge, which can effectively facilitate the training of language modeling while retaining strong visual generation capabilities.
@@
Despite the overall similarity with AR training paradigms, DLMs face several unique challenges due to their diffusion-based formulation. A major issue lies in loss computation efficiency of masked DLMs. In typical masked DLM training, only  50% of tokens (on average) are involved in the loss computation, if uniformly sampling timesteps. This reduces data utilization and may lead to suboptimal gradients, particularly if critical answer tokens are excluded from the loss. To address this, LaViDa [96] proposes a complementary masking strategy: each training sample is duplicated with two disjoint masking patterns, ensuring that all tokens are included in the loss computation at least once. Furthermore, due to the train-inference discrepancy, as illustrated in  [146], the model performs significantly better during training than at inference time. The authors propose a two-step diffusion process and an improved scheduling technique to mitigate this issue.
@@ 3.2 Post-training for Reasoning Capabilities
Exploration of reasoning capabilities is becoming increasingly popular in DLMs as their performance on language tasks improves. Typically, reasoning capabilities are gained through fine-tuning on reasoning datasets. For DLMs, this presents a unique and formidable challenge. Traditional Chain-of-Thought (CoT) methods are based on the sequential nature of AR models to reason step-by-step, but DLMs generate tokens in parallel. The most successful post-training techniques in the AR domain, particularly those based on reinforcement learning (RL) and policy gradient methods, are built upon the ability to efficiently compute the log-probability of a generated sequence. This is straightforward in AR models due to their factorizable, sequential nature. In DLM, where generation is an iterative, non-sequential process, the log-likelihood is intractable, creating a significant technical barrier to applying the mature suite of RL algorithms to AR models. Intuitively, we categorized these works into three main streams, which form the structure of this subsection: (1) Parallelizing the reasoning chain, where CoT in AR models is adapted to DLMs in parallel generation. (2) Adapting policy gradient methods, where variants of popular algorithms like GRPO are introduced to DLMs. (3) Adapting preference optimization methods such as DPO to DLMs.
@@ 3.2.2 Adapting Policy Gradient Methods to DLMs
Score Entropy Policy Optimization (SEPO) [83] introduce RLHF to discrete DLMs, proposing a theoretically grounded framework to fine-tune discrete diffusion models using policy gradient methods and non-differentiable rewards. Operating within the score entropy framework, SEPO adapts modern policy gradient methods like PPO and GRPO by using importance sampling to derive a stable and low-variance gradient estimate. This allows the model’s policy to be iteratively updated to maximize a reward function, making it a general framework for both conditional and unconditional generation. The objective function of SEPO is defined as follows:
@@ 4.1 Parallel Decoding
Parallel decoding naturally aligns with DLMs, leveraging their inherent mask-predict capability to generate multiple tokens simultaneously rather than sequentially. However, naïve parallelization can degrade coherence, motivating a series of adaptive strategies that balance efficiency and quality. Fast-dLLM [44] employs confidence-aware decoding, selectively unmasking tokens whose predicted probabilities exceed a threshold, and realizes up to 27.6× speed-ups without compromising quality. Adaptive Parallel Decoding (APD) [45] modulates the degree of parallelism on-the-fly by consulting a lightweight autoregressive auxiliary model, thus trading throughput for fidelity when necessary. SlowFast Sampling [86] introduces a two-stage schedule. Firstly, a cautious “slow” phase to locate stable tokens, then an aggressive “fast” phase to finalise them in bulk, achieving up to 34× acceleration when combined with caching. SpecDiff [78] further pushes throughput by using a discrete diffusion model as a fully parallel ’drafter’ whose output is quickly verified (and corrected if needed) by a larger autoregressive model, yielding up to 7.2× speed-ups over vanilla AR generation. Finally, the vision–language model Dimple [30] employs confident parallel decoding, dynamically adapting the number of tokens revealed per step and cutting generation iterations by 1.5 – 7×. Collectively, these parallel decoding approaches substantially narrow the latency gap between diffusion and autoregressive models while preserving, and in some cases improving, generation quality.
@@ 4.2 Unmasking/Remasking
State-of-the-art open-source discrete DLMs such as LLaDA [28] and Dream [26] adopt a mask-predict paradigm: at each diffusion step they unmask high-confidence tokens and remask uncertain positions, iteratively refining the sequence. The choice of unmasking/remasking policy, i.e., low-confidence sampling, random selection, or adaptive temperature, therefore dominates both generation quality and convergence speed, making it one of the most critical inference levers. Early work Masked DLM [63] formalised two baselines: random remasking and confidence-ranked remasking, showing that prioritising low-confidence positions yields better quality at no extra cost. Building on this insight, Fast-dLLM [44] introduces confidence-aware parallel decoding: every step it unmasks all positions whose predicted probabilities exceed a global threshold, realizing up to 13× speed-ups while maintaining accuracy. Most recently, ReMDM [46] proposes a principled inference-time remasking sampler that can remask already decoded tokens for further refinement; by scaling the remasking budget, it offers a smooth compute–quality trade-off and closes the quality gap with autoregressive models under fixed compute. Collectively, these adaptive unmasking/remasking strategies substantially boost the efficiency and quality of diffusion language models, and they integrate cleanly with orthogonal accelerators that will be discussed later such as caching and step distillation.
@@ 4.3 Guidance
where $\lambda$ is the guidance scale that balances fidelity to the condition against sample diversity. This simple formulation now underpins most text-to-image systems (e.g., Stable Diffusion [8]) and has been adopted by DLMs for prompt-controlled generation. Subsequent work refines CFG along several axes: dropout-augmented CFG smooths the quality–diversity curve; particle-based guidance blends multiple conditions; and $p2$-weighting rescales the noise term to stabilise high-$\lambda$ sampling. In the text domain, newer schemes extend guidance to structural and semantic constraints. FreeCache [88] couples a lightweight autoregressive verifier with a discrete DLM: the verifier approves (or vetoes) draft tokens before they are committed, simultaneously enforcing coherence and enabling aggressive feature caching. DINGO [89] formulates regular-expression control as a dynamic-programming search over a DFA, guaranteeing constraint satisfaction without altering the model distribution. In other discrete DLMs, guidance can also be applied at each diffusion step, optionally combined with masking/remasking or caching, to steer content (e.g., topic, sentiment) while preserving efficiency. Overall, guidance has become a cornerstone of diffusion inference, offering a lightweight, tunable handle for aligning model outputs with user intent.
@@ 4.4 Efficient Inference
[…]
@@
Feature Cache. Feature caching was first introduced by DeepCache [90], which leverages the strong similarity of intermediate U-Net activations across consecutive diffusion steps to avoid redundant computation. Follow-up work $\Delta$-DiT [91], Learning-to-Cache [92], and FasterCache [93] demonstrate that the same principle transfers cleanly to Transformer-based diffusion models, yielding comparable speed-ups without retraining. With the rise of diffusion language models, dLLM-Cache [47] extends feature caching to text by distinguishing two redundancies: prompt tokens remain almost static throughout denoising, whereas response tokens evolve only sparsely. It therefore pairs a long-interval prompt cache with an adaptive short-interval response cache refreshed only when a lightweight value-similarity test (“V-verify”) detects substantial change, achieving up to $9\times$ end-to-end speed-ups on LLaDA-8B and Dream-7B. Most recently, FreeCache [88] caches the KV/feature projections of already “clean” tokens and refreshes only dynamic positions, pushing acceleration further to $34\times$ while preserving fidelity. Collectively, these advances illustrate that feature caching can bring diffusion language models within striking distance of autoregressive LLMs in inference latency without sacrificing output quality.
@@
Step Distillation. Step distillation is a widely adopted acceleration technique for diffusion models, collapsing the typical thousand-step denoising process into only a few and sometimes even a single sampling steps, thereby drastically reducing inference time. Unlike the training-free methods discussed earlier, it imposes an offline cost: a compact student network must first be trained to mimic the teacher. Early work such as Progressive Distillation [95], followed by ADD [148] and LADD [149], progressively halves the step count or aligns intermediate distributions to preserve fidelity. Di4C [94] extends the framework to discrete diffusion by explicitly distilling inter-token correlations, enabling four to ten steps students that match teacher quality while providing $\sim$2× speed-ups. Most recently, DLM-One [43] employs score-based distillation with adversarial regularisation to train a continuous diffusion language model that generates an entire sequence in a single forward pass, realising up to 500× acceleration with near-teacher quality. Collectively, these works establish step distillation as the principal route toward closing the latency gap between diffusion and autoregressive language models.

### v2

@@ 2.4 Hybrid AR-Diffusion Language Models
Hybrid AR-Diffusion models aim to strike a balance between the full parallelism of non-autoregressive models and the strong causal dependency modeling of autoregressive models. A prominent strategy for hybrid AR-diffusion modeling adopts a block-wise semi-autoregressive generation process. In this setting, the model generates blocks of tokens autoregressively, while the tokens within each block are generated in parallel using a diffusion-like iterative process. Early efforts such as SSD-LM [76] pioneered hybrid approaches by a block-wise continuous diffusion process on simplex representations, AR-DIFFUSION [77] illustrates a multi-level diffusion process and achieves semi-autoregressive by adjusting timestep according token position. Recent representative model BD3-LM [78] further advances this direction on discrete models, demonstrating strong performance compared to pure AR and diffusion models. CtrlDiff [79] improves this paradigm by introducing dynamic block prediction techniques to enhance block-level efficiency and control. SDAR [81] further strengthens this hybrid paradigm by converting a pretrained autoregressive model into a blockwise diffusion model through a lightweight adaptation stage. It preserves AR-level performance while enabling efficient parallel intra-block generation, achieving scalable speedups without sacrificing quality.
@@
Beyond block-based approaches that combine AR and diffusion at sequence level, hybridization can also occur at the architectural level, where some part of the neural network, typically the encoder, diffuses the entire sequence altogether to an intermediate representation, then an autoregressive decoder generates the final sequence [166]. LADIDA [167] is a slightly different approach that diffuses at document level but decodes sentences by an AR decoder. SpecDiff [80] proposes a collaborative speculative decoding framework, where a lightweight diffusion model drafts candidate outputs, which are then validated and finalized by a large AR model. TiDAR [82] proposes a sequence-level hybrid architecture that integrates diffusion-based parallel drafting and autoregressive sampling within a single forward pass through structured causal-bidirectional attention. It effectively unifies the efficiency of diffusion models with the quality of AR decoding, achieving up to $5\times$ throughput improvements while maintaining AR-level performance. SDLM [83] introduces the Next Sequence Prediction (NSP) paradigm, which unifies next-token and next-block prediction to enable adaptive-length generation. By retrofitting pretrained autoregressive models with parallel block training and confidence-based dynamic decoding, SDLM achieves efficient diffusion-style intra-block generation while remaining KV-cache compatible.
@@ 3.1 Pre-training and Supervised Fine-tuning
To accelerate training, particularly for large-scale models, it is common practice to initialize DLMs from pretrained AR language models or image diffusion models [168, 27]. DiffuGPT and DiffuLLaMA [27] try to initialize masked DLMs with open-sourced LLMs from 127M to 7B parameters, found that DLMs can be efficiently adapted from AR models, significantly reducing training time and cost while achieving comparable or even superior performance to their AR counterparts. Building on this insight, Dream-7B is initialized from Qwen 2.5 7B [164], and is reported to outperform both LLaDA-8B and LLaMA3-8B on various benchmarks. Some multimodal DLMs, on the other hand, are initialized from pretrained image diffusion models. D-DiT [84] and Muddit [85] are initialized from pretrained MM-DiT backbones from SD3 [11] and Meissonic [169] respectively. Although these models are not originally designed for text generation, their latent representations contain intrinsic language-aligned knowledge, which can effectively facilitate the training of language modeling while retaining strong visual generation capabilities.
In terms of scaling properties, recent scaling-law analyses [170, 171] reveal that DLMs exhibit distinct compute-data tradeoffs from AR models: they are substantially more data-hungry under compute constraints, yet possess far greater data reuse potential under multi-epoch training, offering a principled foundation for designing optimal DLM training regimes.
@@
Despite the overall similarity with AR training paradigms, DLMs face several unique challenges due to their diffusion-based formulation. A major issue lies in loss computation efficiency of masked DLMs. In typical masked DLM training, only $\sim$50% of tokens (on average) are involved in the loss computation, if timesteps are sampled uniformly. This reduces data utilization and may lead to suboptimal gradients, particularly if critical answer tokens are excluded from the loss. To address this, LaViDa [110] proposes a complementary masking strategy: each training sample is duplicated with two disjoint masking patterns, ensuring that all tokens are included in the loss computation at least once. Furthermore, due to the train-inference discrepancy, as illustrated in  [172], the model performs significantly better during training than at inference time. The authors propose a two-step diffusion process and an improved scheduling technique to mitigate this issue.
@@ 3.2 Post-training for Reasoning Capabilities
Exploration of reasoning capabilities is becoming increasingly popular in DLMs as their performance on language tasks improves. Typically, reasoning capabilities are gained through fine-tuning on reasoning datasets. For DLMs, this presents a unique and formidable challenge. Traditional Chain-of-Thought (CoT) methods are based on the sequential nature of AR models to reason step-by-step, but DLMs generate tokens in parallel. The most successful post-training techniques in the AR domain, particularly those based on reinforcement learning (RL) and policy gradient methods, are built upon the ability to efficiently compute the log-probability of a generated sequence. This is straightforward in AR models due to their factorizable, sequential nature. In DLM, where generation is an iterative, non-sequential process, the log-likelihood is intractable, creating a significant technical barrier to applying the mature suite of RL algorithms developed for AR models to DLMs. Intuitively, we categorize these works into three main streams, which form the structure of this subsection: (1) Parallelizing the reasoning chain, where CoT in AR models is adapted to DLMs in parallel generation. (2) Adapting policy gradient methods, where variants of popular algorithms like GRPO are introduced to DLMs. (3) Adapting preference optimization methods such as DPO to DLMs.
@@ 3.2.1 DoT and DCoLT: Parallelizing the Reasoning Chain
IGPO [91] Policy Gradient (GRPO) Use inpaint ability in DLMs Inpainting-guided sampling; insert partial Discrete
to guide exploration gt reasoning traces when sampling
wd1 [90] Policy Gradient Mitigate computational overhead Weighted likelihood requiring only Discrete
and bias one approximation
SAPO [93] Policy Gradient (GRPO) Learn structured coherent reasoning Process-based reward function Discrete
SPG [92] Policy Gradient Reduce bias in single bound Sandwiched by upper and lower bound; Discrete
methods block-wise masking
BGPO [94] Policy Gradient Reduce memory overhead, scale to Boundary-guided lower bound; Discrete
larger MC sample size constant memory via gradient accumulation
@@ 3.2.2 Adapting Policy Gradient Methods to DLMs
Score Entropy Policy Optimization (SEPO) [88] introduces RLHF to discrete DLMs, proposing a theoretically grounded framework to fine-tune discrete diffusion models using policy gradient methods and non-differentiable rewards. Operating within the score entropy framework, SEPO adapts modern policy gradient methods like PPO and GRPO by using importance sampling to derive a stable and low-variance gradient estimate. This allows the model’s policy to be iteratively updated to maximize a reward function, making it a general framework for both conditional and unconditional generation. The objective function of SEPO is defined as follows:
@@
Sandwiched Policy Gradient (SPG) [92] leverages both an upper and a lower bound of the true log-likelihood to reduce bias in single-sided approximation policy gradient methods for DLMs. The two bounds of likelihood are estimated via Monte Carlo using a block-wise masking strategy to improve training stability. SPG reports to achieve state-of-the-art performance compared with baseline methods on various reasoning benchmarks when applied to LLaDA.
wd1 [90] introduces a novel policy optimization approach that reformulates the objective as a weighted likelihood, requiring only a single approximation for the current parametrized policy likelihood. This formulation reduces bias and improves both stability and training efficiency, outperforming prior diffusion-based RL methods by up to 16% accuracy on reasoning tasks.
IGPO [91] leverages the unique inpainting ability of masked diffusion models to guide exploration during reinforcement learning. By partially injecting ground-truth reasoning traces, IGPO alleviates the zero-advantage problem in group-based RL.
SAPO [93] proposes a step-aware policy optimization scheme that introduces fine-grained process rewards aligned with the latent reasoning hierarchy, mitigating “unstructured refinement” and yielding more interpretable multi-step reasoning traces.
To improve memory efficiency in ELBO-based RL for DLMs, BGPO [94] introduces a boundary-guided lower bound that allows large Monte Carlo sample sizes without increased memory usage, achieving stronger reasoning performance under the same hardware limits.
@@ 4.1 Parallel Decoding
[…]
@@ 4.2 Unmasking/Remasking
State-of-the-art open-source discrete DLMs such as LLaDA [28] and Dream [26] adopt a mask-predict paradigm: at each diffusion step they unmask high-confidence tokens and remask uncertain positions, iteratively refining the sequence. The choice of unmasking/remasking policy, i.e., low-confidence sampling, random selection, or adaptive temperature, therefore dominates both generation quality and convergence speed, making it one of the most critical inference levers. Early work Masked DLM [64] formalized two baselines: random remasking and confidence-ranked remasking, showing that prioritizing low-confidence positions yields better quality at no extra cost. Building on this insight, Fast-dLLM [45] introduces confidence-aware parallel decoding: every step it unmasks all positions whose predicted probabilities exceed a global threshold, realizing up to 13× speedups while maintaining accuracy. Most recently, ReMDM [47] proposes a principled inference-time remasking sampler that can remask already decoded tokens for further refinement; by scaling the remasking budget, it offers a smooth compute–quality trade-off and closes the quality gap with autoregressive models under fixed compute. Collectively, these adaptive unmasking/remasking strategies substantially boost the efficiency and quality of diffusion language models, and they integrate cleanly with orthogonal accelerators that will be discussed later such as caching and step distillation.
@@ 4.3 Guidance
where $\lambda$ is the guidance scale that balances fidelity to the condition against sample diversity. This simple formulation now underpins most text-to-image systems (e.g., Stable Diffusion [8]) and has been adopted by DLMs for prompt-controlled generation. Subsequent work refines CFG along several axes: dropout-augmented CFG smooths the quality–diversity curve; particle-based guidance blends multiple conditions; and $p2$-weighting rescales the noise term to stabilize high-$\lambda$ sampling. In the text domain, newer schemes extend guidance to structural and semantic constraints. FreeCache [100] couples a lightweight autoregressive verifier with a discrete DLM: the verifier approves (or vetoes) draft tokens before they are committed, simultaneously enforcing coherence and enabling aggressive feature caching. DINGO [101] formulates regular-expression control as a dynamic-programming search over a DFA, guaranteeing constraint satisfaction without altering the model distribution. In other discrete DLMs, guidance can also be applied at each diffusion step, optionally combined with masking/remasking or caching, to steer content (e.g., topic, sentiment) while preserving efficiency. Overall, guidance has become a cornerstone of diffusion inference, offering a lightweight, tunable handle for aligning model outputs with user intent.
@@ 4.4 Efficient Inference
Key-Value Cache. The conventional KV cache leverages the strictly autoregressive decoding pattern of LLMs and is therefore ill-suited to the bidirectional, multi-step generation paradigm of DLMs [49]. Recent work, however, shows that carefully redesigning the decoding schedule can recover much of its benefit. Block Diffusion [78] introduces Block Discrete Denoising Diffusion Language Models (BD3-LMs), which decode text autoregressively across coarse blocks while running diffusion within each block; once a block is finished, its keys and values are frozen and reused, enabling variable-length generation and measurable speedups. Fast-dLLM [45] keeps the blockwise view but adds a training-free, approximate DualCache that exploits the near-identity of KV activations across successive diffusion steps for both prefix and suffix tokens, delivering up to $27\times$ end-to-end throughput gains on LLaDA and Dream with $1\%$ accuracy loss. Complementing these block-based schemes, dKV-Cache [49] observes that token representations stabilize only after a position is decoded and therefore deploys a delayed, conditional cache that stores KVs one step later; this design achieves $2$-$10\times$ speedups on the same models with negligible quality drop. d2Cache [103] introduces a fine-grained dual adaptive caching scheme, adaptively refreshes only rapidly changing KV states while reusing stable ones. Elastic-Cache [102] proposes an attention/depth-aware adaptive refresh mechanism that selectively updates deeper layers and reuses stable shallow-layer caches. It performs attention-based drift detection to trigger cache refreshes only when the most-attended tokens exhibit significant changes, yielding up to $45\times$ speedup with minimal quality loss. Together, these results show that semi-autoregressive scheduling and delayed caching provide practical bridges between diffusion’s bidirectional conditioning and Transformer tricks originally devised for autoregression.
@@
Feature Cache. Feature caching was first introduced by DeepCache [104], which leverages the strong similarity of intermediate U-Net activations across consecutive diffusion steps to avoid redundant computation. Follow-up work $\Delta$-DiT [105], Learning-to-Cache [106], and FasterCache [107] demonstrate that the same principle transfers cleanly to Transformer-based diffusion models, yielding comparable speedups without retraining. With the rise of diffusion language models, dLLM-Cache [48] extends feature caching to text by distinguishing two redundancies: prompt tokens remain almost static throughout denoising, whereas response tokens evolve only sparsely. It therefore pairs a long-interval prompt cache with an adaptive short-interval response cache refreshed only when a lightweight value-similarity test (“V-verify”) detects substantial change, achieving up to $9\times$ end-to-end speedups on LLaDA-8B and Dream-7B. Most recently, FreeCache [100] caches the KV/feature projections of already “clean” tokens and refreshes only dynamic positions, pushing acceleration further to $34\times$ while preserving fidelity. Collectively, these advances illustrate that feature caching can bring diffusion language models within striking distance of autoregressive LLMs in inference latency without sacrificing output quality.
@@
Step Distillation. Step distillation is a widely adopted acceleration technique for diffusion models, collapsing the typical thousand-step denoising process into only a few and sometimes even a single sampling steps, thereby drastically reducing inference time. Unlike the training-free methods discussed earlier, it imposes an offline cost: a compact student network must first be trained to mimic the teacher. Early work such as Progressive Distillation [109], followed by ADD [174] and LADD [175], progressively halves the step count or aligns intermediate distributions to preserve fidelity. Di4C [108] extends the framework to discrete diffusion by explicitly distilling inter-token correlations, enabling four to ten steps students that match teacher quality while providing $\sim$2× speedups. Most recently, DLM-One [44] employs score-based distillation with adversarial regularisation to train a continuous diffusion language model that generates an entire sequence in a single forward pass, achieving up to 500× acceleration with near-teacher quality. Collectively, these works establish step distillation as the principal route toward closing the latency gap between diffusion and autoregressive language models.
@@ 5 Multimodal and Unified Approaches
Lavida-O [114] further extends LaViDa into a full-spectrum unified multimodal model capable of both high-quality image generation and fine-grained understanding tasks. It introduces a novel Elastic Mixture-of-Transformers (Elastic-MoT) architecture that decouples the model into a lightweight generation branch and a more powerful understanding branch, enabling scalable training and inference. Lavida-O uniquely supports localized object-level understanding, instruction-based image editing, high-resolution text-to-image synthesis (1024px), and interleaved reasoning and planning within a single unified diffusion framework.
@@
MMaDA-Parallel [115] replaces the sequential reasoning-then-generation pipeline in MMaDA with a fully parallel multimodal diffusion framework, enabling text and images to interact bidirectionally at every denoising step. By jointly generating reasoning traces and visual outputs and further optimizing cross-modal consistency via a trajectory-level Parallel RL (ParaRL) algorithm, MMaDA-Parallel substantially improves semantic alignment and thinking-aware image synthesis performance.